OpenStudy (anonymous):

if i have a function y= t +4x^2 - 10x .. and t = e^x ... how do you calculate dy/dx according to the chain rule ??? can you say that dy/dx = dy/dt*dt/dx ?? i don't understand how it can be done !!

OpenStudy (nenadmatematika):

plug in t=e^x and do derivative of y(x)....=e^x+8x-10

OpenStudy (anonymous):

yes i know .. but i'm asking if you can write according to the chain rule that dy/dx= dy/dt*dt/dx .. because i don't see it